OpenStudy (please_help):

Evaluate the expression. -d⁴= when d= -3 -d⁴= Please simplify the answer and type as integer or decimal.

OpenStudy (anonymous):

-81

OpenStudy (anonymous):

-d^4=3^4 =3*3*3*3 =81..

OpenStudy (please_help):

Thank you!